# Distal Causal Excursion Effects: Modeling Long-Term Effects of Time-Varying Treatments in Micro-Randomized Trials

### v1

@@ Abstract
Keywords: causal inference, distal causal excursion effect, distal outcome, estimating equations, longitudinal data analysis, micro-randomized trials
@@ 1 Introduction
Micro-randomized trials (MRTs) are widely used to develop and optimize digital adaptive interventions (Klasnja et al., 2015; Liao et al., 2016), with applications in health, public policy, education, and information systems (Walton et al., 2018; NeCamp et al., 2020; Liu et al., 2023; Breitwieser et al., 2024; Pieper et al., 2024). In an MRT, each participant is sequentially randomized among treatment options (such as receiving or not receiving a push notification encouraging behavioral change), often hundreds or even thousands of times per participant. These interventions directly target proximal outcomes (i.e., short-term behavioral responses to treatment, such as step count in the next 30 minutes), but ultimately aim to influence distal outcomes (i.e., long-term behaviors and habit formation).
@@
To analyze MRT data, the causal excursion effect (CEE) was introduced to model the effect of time-varying treatments on proximal outcomes (Boruvka et al., 2018; Qian et al., 2021). The CEE, along with the associated estimation methods proposed in those works, has become the standard approach to conduct primary and secondary analyses for MRTs (Klasnja et al., 2018; Nahum-Shani et al., 2021; Bell et al., 2023; Wang et al., 2023; Arévalo Avalos et al., 2024). However, these approaches do not apply to distal outcomes, leaving key scientific questions unanswered. For example, in a physical activity study (Klasnja et al., 2015), do activity prompts that target step count in the short term lead to sustained habit formation? In a smoking cessation study (Battalio et al., 2021), do mindfulness reminders help participants internalize skills to manage cravings after the intervention ends? Current practice relies on behavioral theory to assume the existence of long-term effects; thus, statistical methods that can infer the long-term effects from MRT data are critically needed.
@@
To address this gap, we propose the distal causal excursion effect (DCEE) to quantify the impact of time-varying treatments on a distal outcome measured at the end of an MRT. The DCEE is a contrast between the distal outcomes under two treatment policies that represent excursions (i.e., deviations) from the MRT treatment policy. The DCEE definition implicitly averages over the stochastically assigned treatment at most decision points, thus enabling a parsimonious causal effect model even with a large number of decision points; a related marginalization idea was used in defining CEE for proximal outcomes. This marginalization overcomes a fundamental challenge in MRT analysis: the treatments are randomized at hundreds or even thousands of decision points. Classical causal models, such as marginal structural models and structural nested mean models (Robins, 2000), perform well when the number of treatment occasions is limited. However, in MRTs with many decision points, these models become impractical unless restrictive structural assumptions are imposed—assumptions that are often unrealistic in the behavior change context (Rudolph et al., 2022; Zhang et al., 2022).
@@
To estimate the DCEE, we develop two estimators—one with cross-fitting and one without—both robust to outcome model misspecification. We establish their asymptotic properties and validate their finite-sample performance through simulations. We apply the proposed method to the HeartSteps MRT to assess the long-term effect of activity suggestions on habit formation measured at the end of the study. Our analysis reveals that activity suggestions delivered earlier in the study, despite being farther from the distal outcome, have a larger effect compared to those delivered later. We discuss the implications of these findings for designing future digital interventions.
@@
Section 2 gives the definition of the distal causal excursion effect. Section 3 presents the estimators and the asymptotic theory. Section 4 presents the simulation study. Section 5 presents the real data application. Section 6 concludes with discussion.
@@ 2.1 Notation: MRT with a Distal Outcome
Consider an MRT with $n$ individuals, each in the trial for $T$ decision points, where treatment is randomized. In the following, variables without the subscript $i$ correspond to a generic individual. Let $A_{t}$ denote the binary treatment assignment at decision point $t$, where $A_{t}=1$ indicates treatment and $A_{t}=0$ indicates no treatment. Let $X_{t}$ denote observations recorded between decision points $t-1$ and $t$. The distal outcome, denoted by $Y$, is measured at the end of the study and typically represents the primary health outcome of interest. The observed data trajectory for an individual is given by $O=(X_{1},A_{1},X_{2},A_{2},\ldots,X_{T},A_{T},Y)$. Overbars indicate sequences of variables; for example, $\bar{A}_{t}=(A_{1},A_{2},\ldots,A_{t})$. The history of observed information for an individual up to decision point $t$ (excluding $A_{t}$) is $H_{t}=(X_{1},A_{1},\ldots,X_{t-1},A_{t-1},X_{t})$. At each decision point $t$, $A_{t}$ is assigned according to a randomization probability $p_{t}(a|H_{t}):=P(A_{t}=a\mid H_{t})$ for $a\in\{0,1\}$. Data from individuals are assumed to be independent and identically distributed samples from the distribution $P_{0}$, and expectations are taken with respect to $P_{0}$ unless stated otherwise. The data structure is illustrated in Fig. 1.
@@
In designing the MRT, researchers may deem it unsafe or unethical to deliver push notifications at specific times, such as when a participant is driving. At such decision points, the participant is considered ineligible for randomization, and no treatment will be delivered. Formally, $X_{t}$ includes an indicator $I_{t}$, with $I_{t}=1$ denoting being eligible for randomization at decision point $t$, and $I_{t}=0$ otherwise. If $I_{t}=0$, then $A_{t}=0$ deterministically. In the MRT literature $I_{t}=0$ is also referred to as being unavailable for treatment (Boruvka et al., 2018). We adopt the terminology “ineligible” for clarity.
Let $\mathbb{P}_{n}$ denote the empirical mean over all individuals. For any positive integer $k$, define $[k]:=\{1,2,\ldots,k\}$. The superscript $\star$ is used to indicate quantities associated with the true data-generating distribution $P_{0}$. For a vector $\alpha$ and a vector-valued function $f(\alpha)$, the notation $\partial_{\alpha}f(\alpha):=\partial f(\alpha)/\partial\alpha^{T}$ denotes the Jacobian matrix, where the $(i,j)$-th entry corresponds to the partial derivative of the $i$-th entry of $f$ with respect to the $j$-th entry of $\alpha$.
@@ 2.2 Distal Causal Excursion Effect (DCEE)
To define the causal effect, we use the potential outcomes notation (Rubin, 1974; Robins, 1986). Lowercase letters represent instantiations (non-random values) of the corresponding capital letter random variable. For example, $a_{t}$ is an instantiation of treatment $A_{t}$. For every individual, denote by $X_{t}(\bar{a}_{t-1})$ the $X_{t}$ that would have been observed at decision point $t$ if the individual were assigned a treatment sequence of $\bar{a}_{t-1}$ prior to $t$. The potential outcome of $H_{t}$ under $\bar{a}_{t-1}$ is $H_{t}(\bar{a}_{t-1})=\{X_{1},a_{1},X_{2}(a_{1}),a_{2},\ldots,X_{t}(\bar{a}_{t-1})\}$. The potential outcome of $Y$ under $\bar{a}_{T}$ is $Y(\bar{a}_{T})$.
We use the term policy to refer to any decision rule (static or dynamic) of assigning $A_{t}$ given the history $H_{t}$ for $t\in[T]$ (Murphy, 2003). The MRT policy, denoted by $D:=(d_{1},\ldots,d_{T})$, is the rule in the MRT to stochastically assign $A_{t}$ for every $t\in[T]$. Specifically, $d_{t}(\cdot)$ is a stochastic mapping that maps $H_{t}$ to $\{0,1\}$, such that $d_{t}(H_{t})=a$ with probability $P(A_{t}=a\mid H_{t})$ for $a\in\{0,1\}$. For each $t$, we consider two alternative decision rules for assigning $A_{t}$ that will be contrasted in defining the causal effect: $d_{t}^{1}$ is the decision rule that always assigns $A_{t}=1$ unless $I_{t}=0$ (in which case $A_{t}=0$), and $d_{t}^{0}$ always assigns $A_{t}=0$. In other words, $d_{t}^{1}(H_{t})=I_{t}$, and $d_{t}^{0}(H_{t})=0$. For $a\in\{0,1\}$, let $D_{d_{t}=d_{t}^{a}}$ denote the policy obtained by replacing $d_{t}$ with $d_{t}^{a}$ in the MRT policy $D$:
@@
$D_{d_{t}=d_{t}^{1}}$ and $D_{d_{t}=d_{t}^{0}}$ are two excursions (Qian et al., 2021; Guo et al., 2021) from the MRT policy $D$ in that they deviate from $D$ at decision point $t$ to always assign $A_{t}=1$ (while respecting the eligibility constraint) or always assign $A_{t}=0$, respectively.
Let $S_{t}$ denote a subset of $H_{t}$, which are the moderators of interest based on the scientific question. For example, one can set $S_{t}=\emptyset$ for a fully marginal effect that averages over all moderators, $S_{t}=A_{t-1}$ for effect moderation by previous intervention, or $S_{t}=X_{t}$ for effect moderation by current covariates. Note that $S_{t}$ can be time-varying. We consider the following distal causal excursion effect (DCEE) of $A_{t}$ on $Y$:
@@
$\tau(t,s)$ captures the difference in the potential outcomes of $Y$ contrasting two excursion policies, $D_{d_{t}=d_{t}^{1}}$ and $D_{d_{t}=d_{t}^{0}}$, conditional on effect modifiers $S_{t}$. In other words, $\tau(t,s)$ contrasts what would have happened if a participant followed the MRT policy throughout the study except for decision point $t$ where they received $A_{t}=I_{t}$ versus $A_{t}=0$. Technically, the notation $\tau(t,s)$ should further depend on the choice of $S_{t}\subset H_{t}$, but we omit this notational dependence to avoid unnecessary complexity. The choice of $S_{t}$ will be clear from the context.
Below we present several examples with different data-generating distributions to illustrate the DCEE and highlight its scientific relevance for digital interventions. In all examples, we assume that a larger $Y$ is desirable, the error term $\epsilon$ is exogenous with mean 0, and we set $S_{t}=\emptyset$ to focus on fully marginal effects $\tau(t)$. A time-varying variable is exogenous if it is independent of its own history and the history of other variables. More general versions of these examples, along with detailed derivations of the DCEE, are provided in the Supplementary Material A. Importantly, the estimators proposed in Section 3 are valid as long as the assumptions in Theorem 2 hold; they do not require knowledge of most aspects of the data-generating process, such as how $Y$ depends on $A_{t}$ and $X_{t}$ or how $A_{t}$ influences $X_{t+1}$ and future eligibility $I_{t+1}$.
@@
Suppose that for all $t\in[T]$, the covariate $X_{t}$ is exogenous, all individuals are always eligible ($I_{t}\equiv 1$), and treatment $A_{t}\sim\text{Bern}(p)$ is exogenous. Suppose $Y=g(\bar{X}_{T})+\sum_{t=1}^{T}A_{t}(\alpha_{t}+\beta_{t}X_{t})+\epsilon$ for some unspecified $g$. Since we set $S_{t}=\emptyset$, the DCEE $\tau(t)$ represents the fully marginal effect of $A_{t}$, given by $\tau(t)=\alpha_{t}+\beta_{t}~{}\mathbb{E}(X_{t})$. This illustrates how the DCEE naturally marginalizes over effect modifiers not explicitly included in $S_{t}$.
@@
Consider the same data-generating process as Example 1, but instead suppose that there is a negative interaction between consecutive treatments: $Y=g(\bar{X}_{T})+\sum_{t=1}^{T}\beta_{t}A_{t}-\sum_{t=1}^{T-1}\alpha_{t}A_{t}A_{t+1}+\epsilon$ with $\alpha_{t}\geq 0$. Here, treatment effects diminish when treatments are delivered in succession, capturing user burden. The average-over-time effect is $\frac{1}{T}\sum_{t=1}^{T}\tau(t)=\frac{1}{T}\sum_{t=1}^{T}\{\beta_{t}-(\alpha_{t}+\alpha_{t-1})p\}$, where $\alpha_{0}:=0$. The term $-(\alpha_{t}+\alpha_{t-1})p$ reflects the deterioration of treatment effects due to user burden, which increases with the treatment probability $p$. Thus, the DCEE automatically accounts for user burden, and as a result different MRTs with different $p$ will yield different DCEE.
@@
Consider a two-timepoint setting ($T=2$), where $X_{1}$ is exogenous, all individuals are always eligible ($I_{1}=I_{2}\equiv 1$), treatments $A_{1},A_{2}\sim\text{Bern}(p)$ are exogenous, and the treatment $A_{1}$ influences the time-varying covariate $X_{2}$ through $X_{2}\mid A_{1}\sim\text{Bern}(\rho_{0}+\rho_{1}A_{1})$. Suppose $Y=\gamma_{0}+\gamma_{1}X_{1}+\gamma_{2}X_{2}+\beta_{1}A_{1}+\beta_{2}A_{2}+\epsilon$. In this case, the DCEE at $t=1$ is $\tau(1)=\beta_{1}+\gamma_{2}\rho_{1}$. This expression captures both the direct effect of $A_{1}$ on $Y$ ($\beta_{1}$) and the indirect effect mediated through $X_{2}$ ($\gamma_{2}\rho_{1}$), demonstrating how the DCEE naturally combines mediation pathways in time-varying interventions.
@@
Consider a two-timepoint setting ($T=2$), where all individuals are eligible at $t=1$ ($I_{1}\equiv 1$), treatment at $t=1$ influences future eligibility: $I_{2}\mid A_{1}\sim\text{Bern}(\rho_{0}-\rho_{1}A_{1})$, and the treatment assignments follow $A_{1}\sim\text{Bern}(p)$ and $A_{2}\mid I_{2}=1\sim\text{Bern}(p)$. Suppose $Y=\beta_{0}+\beta_{1}A_{1}+\beta_{2}A_{2}+\epsilon$. Assume $\beta_{1},\beta_{2},\rho_{0},\rho_{1}>0$. In this case, the DCEE at $t=1$ is $\tau(1)=\beta_{1}-\beta_{2}p\rho_{1}$. Here, the term $-\beta_{2}p\rho_{1}$ captures the reduced effect of $A_{1}$ on $Y$ due to its negative impact on future eligibility. This demonstrates how the DCEE accounts for delayed effects of treatment via eligibility constraints, a key consideration in MRTs as eligibility may depend on past treatment history.
@@
The DCEE differs from the causal excursion effect (CEE) (Boruvka et al., 2018; Qian et al., 2021) in several key ways: (i) Outcome focus: DCEE focuses on the end-of-study distal outcome, whereas CEE focuses on short-term proximal outcomes. (ii) Policy comparison: DCEE contrasts treatment policies on the entire sequence $A_{1},\ldots,A_{T}$, while CEE contrasts policies up to time $t$ (or up to $A_{t+\Delta-1}$ for some prespecified $\Delta\geq 1$). (iii) Handling of eligibility: DCEE marginalizes over the eligibility indicator and contrasts policies that respect the eligibility constraints, whereas CEE conditions on a decision point being eligible. The distinction in eligibility handling (iii) is particularly important: DCEE accounts for the long-term consequences of treatment on future eligibility, meaning that a treatment that is highly effective but substantially reduces future eligibility may lead to an attenuated DCEE (see Example 4). This issue does not arise in CEE, as it focuses only on short-term proximal outcomes. ∎
@@
The DCEE also differs from other causal estimands for time-varying treatments, such as those considered in marginal structural model (MSM) or the structural nested mean model (SNMM) (Robins, 2000). In an MRT setting, a MSM would model the expected $Y$ under fixed treatment policies $(a_{1},\ldots,a_{T})$, and a SNMM would model the effect of a treatment blip on $Y$ (i.e., contrasting $A_{t}=1$ versus 00) after removing all future treatments (i.e., after setting $A_{t+1}=\ldots=A_{T}=0$). Without strong structural assumptions such as the Markovian property which are likely infeasible in behavior change settings, both approaches will introduce more parameters than can be reliably estimated from MRT data. In contrast, the DCEE compares stochastic treatment policies that deviate from the MRT policy at only one decision point, allowing for parsimonious modeling while preserving interpretability. This idea is related to the longitudinal modified treatment policies framework (Díaz et al., 2023), but DCEE uniquely accounts for effect modification by time-varying covariates and eligibility constraints, which are central to digital intervention studies. ∎

### v2

@@ Abstract
Keywords: causal excursion effect; distal outcome; micro-randomized trial; mobile health; time-varying treatment
@@ 1 Introduction
Micro-randomized trials (MRTs) are widely used to develop and optimize digital adaptive interventions (Klasnja et al., 2015; Liao et al., 2016), with applications in health, education, and information systems (Walton et al., 2018; Breitwieser et al., 2024; Pieper et al., 2024). In an MRT, each participant is sequentially randomized among treatment options (e.g., receiving or not receiving a push notification encouraging behavioral change). These interventions directly target proximal outcomes (short-term behavioral responses, such as step count in the next 30 minutes), but ultimately aim to influence distal outcomes (long-term behaviors and habit formation).
@@
To analyze MRT data, the causal excursion effect (CEE) was introduced to model the effect of time-varying treatments on proximal outcomes (Boruvka et al., 2018). It has become the standard for primary and secondary analyses in MRTs (Klasnja et al., 2019; Liu et al., 2023). However, this estimand does not address distal outcomes, leaving key scientific questions unanswered. For example, in a physical activity study (Klasnja et al., 2015), do activity prompts lead to sustained habit formation? Current practice relies on behavioral theory to assume long-term effects, highlighting the need for statistical methods to infer them from MRT data.
@@
To address this critical need, we propose the distal causal excursion effect (DCEE) to quantify the impact of time-varying treatments on a distal outcome measured at the end of an MRT. The DCEE contrasts distal outcomes under two treatment policies representing excursions (i.e., deviations) from the MRT policy. By marginalizing over the stochastic treatment assignments at most decision points, the DCEE enables a parsimonious causal effect model even with many decision points; a similar idea was used in CEE for proximal outcomes. This marginalization addresses a key challenge in MRT analysis: treatments are randomized at hundreds or thousands of decision points. Classical models such as marginal structural models and structural nested models (Robins, 1999a) work well with a few treatment occasions but become impractical in MRTs with many decision points unless restrictive and often unrealistic assumptions are imposed (Rudolph et al., 2022).
@@
To estimate the DCEE, we develop two estimators—one with cross-fitting and one without—both robust to outcome model misspecification. We establish their asymptotic properties and validate finite-sample performance through simulations. We apply the method to the HeartSteps MRT to assess the long-term effect of activity suggestions on habit formation at study end. Our analysis reveals that suggestions delivered earlier, despite being farther from the distal outcome, have a larger effect than those delivered later. We discuss the implications for designing future digital interventions.
@@
Section 2 defines the DCEE. Section 3 presents the estimators and the asymptotic theory. Section 4 presents the simulation study. Section 5 presents the real data application. Section 6 concludes with discussion.
@@ 2.1 Notation: MRT with a Distal Outcome
Consider an MRT with $n$ individuals, each in the trial for $T$ decision points where treatment is randomized. Variables without subscript $i$ refer to a generic individual. Let $A_{t}$ denote the binary treatment at decision point $t$: $A_{t}=1$ for treatment, $A_{t}=0$ for no treatment. Let $X_{t}$ denote observations recorded between decision points t1t-1 and $t$. The distal outcome $Y$, typically the primary health outcome, is measured at study end. The observed data trajectory for one individual is $O=(X_{1},A_{1},\ldots,X_{T},A_{T},Y)$. Overbars indicate sequences of variables; for example, $\bar{A}_{t}=(A_{1},\ldots,A_{t})$. The history up to $t$ (excluding $A_{t}$) is $H_{t}=(X_{1},A_{1},\ldots,X_{t-1},A_{t-1},X_{t})$. At each $t$, $A_{t}$ is assigned with randomization probability $p_{t}(a|H_{t}):=P(A_{t}=a\mid H_{t})$ for $a\in\{0,1\}$. Data from individuals are assumed independent and identically distributed samples from the distribution $P_{0}$, and expectations are taken under $P_{0}$ unless stated otherwise. Fig. 1 illustrates the data structure.
@@
In designing the MRT, researchers may deem it unsafe or unethical to deliver push notifications at certain times, such as when a participant is driving. At these decision points, the participant is considered ineligible for randomization, and no treatment is delivered. Formally, $X_{t}$ includes an indicator $I_{t}$, with $I_{t}=1$ denoting being eligible for randomization at $t$, and $I_{t}=0$ otherwise. If $I_{t}=0$, then $A_{t}=0$ deterministically. In the MRT literature, $I_{t}=0$ is also called being unavailable for treatment (Boruvka et al., 2018); here, we use “ineligible” for clarity. For simplicity, we sometimes explicitly write $I_{t}$ together with $H_{t}$ in conditional expectations, even though $I_{t}$ is part of $H_{t}$: for instance, $\mathbb{E}(\cdot|H_{t},I_{t}=1)$ means $\mathbb{E}(\cdot|H_{t}\setminus\{I_{t}\},I_{t}=1)$.
Let $\mathbb{P}_{n}$ denote the empirical mean over everyone. For any positive integer $k$, define $k]:=\{1,2,\ldots,k\}$. The superscript $\star$ is used to indicate quantities associated with the true data-generating distribution $P_{0}$. For a vector $\alpha$ and a vector-valued function $f(\alpha)$, the notation $\partial_{\alpha}f(\alpha):=\partial f(\alpha)/\partial\alpha^{T}$ denotes the Jacobian matrix, where the $(i,j)$-th entry corresponds to the partial derivative of the $i$-th entry of $f$ with respect to the $j$-th entry of $\alpha$.
@@ 2.2 Distal Causal Excursion Effect (DCEE)
To define the causal effect, we use the potential outcomes notation (Rubin, 1974; Robins, 1986). Lowercase letters denote instantiations (non-random values) of corresponding random variables; for example, $a_{t}$ is an instantiation of $A_{t}$. For each individual, let $X_{t}(\bar{a}_{t-1})$ be the value of $X_{t}$ that would have been observed at $t$ if the individual were assigned treatment sequence $\bar{a}_{t-1}$. The potential outcome of $H_{t}$ under $\bar{a}_{t-1}$ is $H_{t}(\bar{a}_{t-1})=\{X_{1},a_{1},X_{2}(a_{1}),a_{2},\ldots,X_{t}(\bar{a}_{t-1})\}$. The potential outcome of $Y$ under $\bar{a}_{T}$ is $Y(\bar{a}_{T})$.
We use the term policy to refer to any decision rule (static or dynamic) for assigning $A_{t}$ given the history $H_{t}$ for $t\in[T]$ (Murphy, 2003). The MRT policy, denoted by $D:=(d_{1},\ldots,d_{T})$, is the rule for stochastically assign $A_{t}$ at each $t$ in the MRT. Specifically, $d_{t}(\cdot)$ is a stochastic mapping from $H_{t}$ to $\{0,1\}$, where $d_{t}(H_{t})=a$ with probability $P(A_{t}=a\mid H_{t})$ for $a\in\{0,1\}$. For each $t$, we consider two alternative decision rules for defining the causal effect: $d_{t}^{1}$, which always assigns $A_{t}=1$ unless $I_{t}=0$ (in which case $A_{t}=0$), and $d_{t}^{0}$, which always assigns $A_{t}=0$. In other words, $d_{t}^{1}(H_{t})=I_{t}$, and $d_{t}^{0}(H_{t})=0$. For $a\in\{0,1\}$, let $D_{d_{t}=d_{t}^{a}}$ denote the policy obtained by replacing $d_{t}$ with $d_{t}^{a}$ in the MRT policy $D$:
@@
$D_{d_{t}=d_{t}^{1}}$ and $D_{d_{t}=d_{t}^{0}}$ are two excursions (Guo et al., 2021) from the MRT policy $D$: they deviate from $D$ at decision point $t$ to always assign $A_{t}=1$ (while respecting eligibility) or always assign $A_{t}=0$, respectively.
Let $S_{t}$ denote a subset of $H_{t}$, representing moderators of interest. For example, setting $S_{t}=\emptyset$ defines a fully marginal effect averaging over all moderators; setting $S_{t}=A_{t-1}$ captures effect moderation by previous intervention; and setting $S_{t}=X_{t}$ captures effect moderation by current covariates. Note that $S_{t}$ can be time-varying. We define the distal causal excursion effect (DCEE) of $A_{t}$ on $Y$ as
@@
$\tau(t,s)$ captures the difference in the potential outcomes of $Y$ under two excursion policies, $D_{d_{t}=d_{t}^{1}}$ and $D_{d_{t}=d_{t}^{0}}$, conditional on $S_{t}$. In other words, $\tau(t,s)$ contrasts what would have happened if a participant followed the MRT policy throughout the study except at decision point $t$, where they received $A_{t}=I_{t}$ versus $A_{t}=0$. The way $I_{t}$ enters the definition here differs from its role in defining (proximal) causal excursion effects (Boruvka et al., 2018); we elaborate on this distinction further in Remark 1.
Below we present several examples with different data-generating distributions to illustrate the DCEE and highlight its scientific relevance for digital interventions. In all examples, we assume larger $Y$ is desirable, the error term $\epsilon$ is exogenous with mean 0, and set $S_{t}=\emptyset$ to focus on fully marginal effects $\tau(t)$. A time-varying variable is exogenous if it is independent of its own history and the history of other variables. More general versions of these examples, along with detailed derivations, are in Supplementary Material A. Importantly, the estimators proposed in Section 3 remain valid as long as the assumptions in Theorem 3.1 hold; they do not require knowledge of most aspects of the data-generating process, such as how $Y$ depends on $A_{t}$ and $X_{t}$ or how $A_{t}$ influences $X_{t+1}$ and future eligibility $I_{t+1}$.
@@
Suppose that for all $t\in[T]$, the covariate $X_{t}$ is exogenous, all individuals are always eligible ($I_{t}\equiv 1$), and treatment $A_{t}\sim\text{Bern}(p)$ is exogenous. Suppose $Y=g(\bar{X}_{T})+\sum_{t=1}^{T}A_{t}(\alpha_{t}+\beta_{t}X_{t})+\epsilon$ for some unspecified $g$ and $E(\epsilon)=0$. Since we set $S_{t}=\emptyset$, the DCEE $\tau(t)$ represents the fully marginal effect of $A_{t}$, given by $\tau(t)=\alpha_{t}+\beta_{t}~\mathbb{E}(X_{t})$. This illustrates how the DCEE naturally marginalizes over effect modifiers not explicitly included in $S_{t}$.
@@
Consider the same data-generating process as Example 1, but instead suppose that there is a negative interaction between consecutive treatments: $Y=g(\bar{X}_{T})+\sum_{t=1}^{T}\beta_{t}A_{t}-\sum_{t=1}^{T-1}\alpha_{t}A_{t}A_{t+1}+\epsilon$ with $\alpha_{t}\geq 0$. Here, treatment effects diminish when treatments are delivered in succession, capturing user burden. The average-over-time effect is $\frac{1}{T}\sum_{t=1}^{T}\tau(t)=\frac{1}{T}\sum_{t=1}^{T}\{\beta_{t}-(\alpha_{t}+\alpha_{t-1})p\}$, where $\alpha_{0}:=0$. The term $(\alpha_{t}+\alpha_{t-1})p$ reflects the deterioration of treatment effects due to user burden, which increases with the treatment probability $p$. Thus, the DCEE automatically accounts for user burden, and as a result different MRTs with different $p$ will yield different DCEE.
@@
Consider a two-timepoint setting ($T=2$), where $X_{1}$ is exogenous, all individuals are always eligible ($I_{1}=I_{2}\equiv 1$), treatments $A_{1},A_{2}\sim\text{Bern}(p)$ are exogenous, and the treatment $A_{1}$ influences the time-varying covariate $X_{2}$ through $X_{2}\mid A_{1}\sim\text{Bern}(\rho_{0}+\rho_{1}A_{1})$. Suppose $Y=\gamma_{0}+\gamma_{1}X_{1}+\gamma_{2}X_{2}+\beta_{1}A_{1}+\beta_{2}A_{2}+\epsilon$. In this case, the DCEE at $t=1$ is $\tau(1)=\beta_{1}+\gamma_{2}\rho_{1}$. This expression captures both the direct effect of $A_{1}$ on $Y$ ($\beta_{1}$) and the indirect effect mediated through $X_{2}$ ($\gamma_{2}\rho_{1}$).
@@
Consider a two-timepoint setting ($T=2$), where all individuals are eligible at $t=1$ ($I_{1}\equiv 1$), treatment at $t=1$ influences future eligibility: $I_{2}\mid A_{1}\sim\text{Bern}(\rho_{0}-\rho_{1}A_{1})$, and the treatment assignments follow $A_{1}\sim\text{Bern}(p)$ and $A_{2}\mid I_{2}=1\sim\text{Bern}(p)$. Suppose $Y=\beta_{0}+\beta_{1}A_{1}+\beta_{2}A_{2}+\epsilon$. Assume $\beta_{1},\beta_{2},\rho_{0},\rho_{1}>0$. In this case, the DCEE at $t=1$ is $\tau(1)=\beta_{1}-\beta_{2}p\rho_{1}$. Here, the term $\beta_{2}p\rho_{1}$ captures the reduced effect of $A_{1}$ on $Y$ due to its negative impact on future eligibility. This demonstrates how the DCEE accounts for delayed effects of treatment via eligibility constraints, a key consideration in MRTs as eligibility may depend on past treatment history.
@@
The DCEE differs from the causal excursion effect (CEE) (Boruvka et al., 2018) in several key ways: (i) Outcome focus: DCEE focuses on the end-of-study distal outcome, whereas CEE focuses on short-term proximal outcomes. (ii) Policy comparison: DCEE contrasts treatment policies on the entire sequence $A_{1},\ldots,A_{T}$, while CEE contrasts policies up to time $t$ (or up to $A_{t+\Delta-1}$ for some prespecified $\Delta\geq 1$). (iii) Eligibility handling: DCEE marginalizes over eligibility and compares policies that respect eligibility constraints, whereas CEE conditions on eligibility at a given decision point. This last distinction is especially important: DCEE accounts for the long-term consequences of treatment on future eligibility. Thus, a treatment that is effective but substantially reduces future eligibility may yield an attenuated DCEE (see Example 4). This issue does not arise with CEE, which focuses only on short-term outcomes. Consequently, one cannot simply repurpose CEE estimation methods (e.g., weighted and centered least squares (Boruvka et al., 2018)) to estimate the DCEE by treating the distal outcome as a proximal outcome at each decision point. Such an approach would fail to account for eligibility probability, thus resulting in biased DCEE estimates. This is illustrated in our simulation results (Section 4).
@@
The DCEE differs from marginal structural models (MSMs) and structural nested mean models (SNMMs) (Robins, 1999a). In an MRT, a MSM models $Y$ under fixed policies, and a SNMM models a treatment blip effect assuming future treatments are set to zero. Without strong (often unrealistic) assumptions, both require more parameters than MRT data can support. In contrast, DCEE compares policies that deviate at a single decision point from the MRT policy, thus enabling parsimonious, interpretable modeling. ∎
Remark 3 (Differences between DCEE and dynamic treatment regime).
The DCEE differs from dynamic treatment regimes (DTRs) (Murphy, 2003; Robins, 2004) in several aspects. First, DTR approaches contrast entire sequences of decision rules (full dynamic policies), whereas DCEE compares “excursion” policies that only differ at a single decision point and match the reference policy elsewhere. Second, the DTR framework aims to optimize distal outcomes by identifying an optimal policy, while DCEE focuses on evaluating the long-term effect of a local treatment decision without optimizing across sequences. Third, DTR is most suitable for designs with few decision points (e.g., sequential multiple assignment randomized trials (Murphy, 2005)), while DCEE is suitable for MRTs with many decision points. These distinctions reflect a bias-variance trade-off: DCEE defines a more parsimonious estimand suited for intensive longitudinal settings.